OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 2) and a directrix at y = -2.

OpenStudy (anonymous):

y^2 = 2x y = 1/2x^2 y^2 = 8x y = 1/8x^2